OpenStudy (anonymous):

The burning of fossil fuels (oil, coal, natural gas) releases carbon dioxide into the atmosphere. Carbon dioxide is known as a greenhouse gas because it tends to trap heat in the atmosphere, causing global warming. In 1995, the United States emitted about 1.4 billion tons of carbon (as carbon dioxide) into the atmosphere. China was second largest contributor, emitting 850 million tons. US emissions were increasing at 1.3% a year, while Chinese emissions were increasing at 4% a year. Assuming these growth rates continued determine the year when China will exceed US in carbon emissions.

OpenStudy (jchick):

US emissions = 1400(1+0.013)ttons Chinese emissions = 850(1+0.04)ttons where t is the time in years from 1995. To find the time in years from 1995 when US and Chinese emissions are equal, solve the following equation for t 1400×1.013t=850×1.04t Can you do this?